# The hereditariness problem for the Černý conjecture

### v1

@@ 4 The lattice of congruences of an automaton
This section is dedicated to the study of the lattice of congruences of a given DFA $\mathcal{A}=(Q,\Sigma,\delta)$. Most of the results presented here are considered folklore, but for completeness—and to make this paper self-contained—we include them with full details. We begin by studying the congruence generated by a fixed pair $(p,q)\in Q\times Q$. For any pair of states $p,q\in Q$, we can define the congruence generated by $\{p,q\}$, denoted by $\langle\{p,q\}\rangle$, as follows:
$\langle\{p,q\}\rangle:=\bigcap\left\{\sigma\in\text{Cong}(\mathcal{A})\ \middle|\ (p,q)\in\sigma\right\}.$
This is well-defined because the set on the right-hand side is nonempty (since the universal relation $\nabla_{\mathcal{A}}\in\text{Cong}(\mathcal{A})$ contains all pairs, in particular $(p,q)$), and the intersection of a family of congruences is still a congruence. Moreover, the construction is symmetric in $p$ and $q$: had we chosen $(q,p)$ instead of $(p,q)$, we would obtain the same congruence. This justifies the use of set-theoretic brackets $\{p,q\}$ rather than ordered pairs. This definition extends naturally to any nonempty subset $S\subseteq\binom{Q}{2}$, where
$\binom{Q}{2}:=\left\{X\subseteq Q\ \middle|\ |X|=2\right\}.$
We define:
$\langle S\rangle:=\bigcap\left\{\sigma\in\text{Cong}(\mathcal{A})\ \middle|\ S\subseteq\sigma\right\}.$
We now state the following lemma.
@@
Let $p,q\in Q$, and define
$S=\{p,q\}\cdot\Sigma^{*}:=\left\{\{t,s\}\ \middle|\ \exists u\in\Sigma^{*}\text{ such that }\{p,q\}\cdot u=\{t,s\}\right\}.$
Then, $\langle\{p,q\}\rangle=\langle S\rangle$.
@@
Without loss of generality, we may treat $S$ as the symmetric set of ordered pairs generated by it, so that $S\subseteq Q\times Q$. Let $\rho$ be any congruence such that $(p,q)\in\rho$. Then, for every $u\in\Sigma^{*}$, it follows that $(p\cdot u,q\cdot u)\in\rho$. In particular, $(p,q)\cdot u=(p\cdot u,q\cdot u)\in\rho$, so $S\subseteq\rho$. Hence, $\rho$ contains $S$, and thus $\langle S\rangle\subseteq\rho$. Since this holds for all such $\rho$, we get:
$\langle\{p,q\}\rangle\subseteq\langle S\rangle.$
Conversely, observe that $\langle S\rangle\subseteq\bigcap\left\{\langle\{s,t\}\rangle\ \middle|\ \{s,t\}\in S\right\}$, and since $(p,q)\in S$, this intersection contains $\langle\{p,q\}\rangle$, which gives the reverse inclusion. Therefore, the two congruences coincide. ∎
We now describe the congruence generated by a pair $p,q\in Q$ in a more concrete way. With a slight abuse of notation, we continue to denote by $S$ the symmetric binary relation consisting of all pairs $(s,t)$ such that $\{s,t\}\in S:=\{p,q\}\cdot\Sigma^{*}$.
@@
With the above notation, $\langle\{p,q\}\rangle$ is the equivalence closure of $S$.
Proof.
Let $\sigma$ be the equivalence closure of $S$. Then $\sigma$ is an equivalence relation by construction. We claim that $\sigma$ is also a congruence. Let $(t,s)\in\sigma$ with $t\neq s$ (the case $t=s$ is trivial), and let $u\in\Sigma^{*}$. Since $\sigma$ is the equivalence closure of $S$, there exists a sequence of states $p_{1},\ldots,p_{k}\in Q$ and words $v_{1},v_{2}\in\Sigma^{*}$ such that:
$t=p\cdot v_{1}\ \mathrel{S}\ p_{1}\ \mathrel{S}\ \cdots\ \mathrel{S}\ p_{k}\ \mathrel{S}\ q\cdot v_{2}=s.$
Applying the action of $u$ to each state in this sequence, and using the fact that $S$ is closed under right action (by definition), we obtain:
$t\cdot u=p\cdot v_{1}u\ \mathrel{S}\ p_{1}\cdot u\ \mathrel{S}\ \cdots\ \mathrel{S}\ p_{k}\cdot u\ \mathrel{S}\ q\cdot v_{2}u=s\cdot u.$
Hence, $(t\cdot u,s\cdot u)\in\sigma$, showing that $\sigma$ is stable under the transition function and thus a congruence. By construction, $\sigma$ is the smallest equivalence relation containing $S$, and from Lemma 4.1, we know that $\langle\{p,q\}\rangle=\langle S\rangle$. Therefore, we conclude that $\sigma=\langle\{p,q\}\rangle$. ∎
In what follows, $\text{Cong}(\mathcal{A})$ will always be regarded as a poset with respect to the inclusion. It may also be regarded as a lattice where for every $\sigma,\tau\in\text{Cong}(\mathcal{A})$ we define $\sigma\wedge\tau:=\sigma\cap\tau$ and, since $\text{Cong}(\mathcal{A})$ is finite, we define as usual
$\sigma\vee\tau:=\cap\{\rho\in\text{Cong}(\mathcal{A})\ |\ \rho\supseteq\sigma,\tau\}$
which is clearly a congruence since the intersection of congruences is a congruence. Note that the maximum is given by $1:=\nabla_{\mathcal{A}}=Q\times Q$, and the minimum by $0:=\Delta_{\mathcal{A}}=\{(q,q):q\in Q\}$. The following lemma easily follows from the definition of atom of a lattice.
Lemma 4.3.
Let $\text{Cong}(\mathcal{A})$ seen as a lattice. Then, for every atom $\sigma\in\text{Cong}(\mathcal{A})$ we have $\sigma=\langle\{p,q\}\rangle$ for some $p,q\in Q$.
We can rephrase the previous lemma as follows: for any atom $\sigma\in\text{Cong}(\mathcal{A})$ and $(p,q)\in\sigma$ such that $p\neq q$, we have $\sigma=\langle\{p,q\}\rangle$. We now exhibit a simple algorithm that allows the computation of all the atoms which probably belongs to the folklore. We start by introducing the following preorder on $\binom{Q}{2}$:
$\{p,q\}\preceq\{t,s\}\iff\exists u\in\Sigma^{*},\ \{p,q\}\cdot u=\{t,s\}$
@@
Lemma 4.4.
Let $\{p,q\},\{t,s\}\in\binom{Q}{2}$ such that $\{p,q\}\sim\{t,s\}$. Then, $\langle\{p,q\}\rangle=\langle\{t,s\}\rangle$.
Proof.
By definition, $\{p,q\}\sim\{t,s\}$ implies that there are $u,v\in\Sigma^{*}$ such that $\{p,q\}\cdot u=\{t,s\}$ and $\{t,s\}\cdot v=\{p,q\}$, that is $\{p,q\}\cdot\Sigma^{*}=\{t,s\}\cdot\Sigma^{*}$. Hence, by Lemma 4.1, this implies $\langle\{p,q\}\rangle=\langle\{t,s\}\rangle\rangle$. ∎
Proposition 4.5.
[…]
Proof.
Let $\{p,q\}\in\binom{Q}{2}$ be such that $\{p,q\}]_{\sim}$ is maximal and let $\tau\in\text{Cong}(\mathcal{A})\setminus\{\Delta_{\mathcal{A}}\}$ be an atom such that $\tau\subseteq\sigma=\langle(p,q)\rangle$. By Lemma 4.3, we know that there exists $\{t,s\}\in\binom{Q}{2}$ such that $\tau=\langle\{t,s\}\rangle$. But since $\tau\subseteq\sigma$, we have that $(t,s)\in\sigma$ which by Lemma 4.2 implies the existence $\{t,p_{1}\},\{p_{2},s\}\in\{p,q\}\cdot\Sigma^{*}$ with $p_{1}\sigma p_{2}$. Therefore, we have $\{p,q\}\preceq\{t,p_{1}\},\{p_{k},s\}$. By the maximality of $\{p,q\}]_{\sim}$, there are $u_{1},u_{2}\in\Sigma^{*}$ such that $\{t,p_{1}\}\cdot u_{1}=\{p,q\},\ \{p_{k},s\}\cdot u_{2}=\{p,q\}$, that is, $(p,q)\in\tau$. Hence, $\langle\{p,q\}\rangle\subseteq\tau$ and thus $\tau=\sigma$. Conversely, assume that $\langle\{p,q\}\rangle$ is an atom of $\text{Cong}(\mathcal{A})$ and $\{p,q\}]_{\sim}\preceq[\{t,s\}]_{\sim}$. Then, $\{t,s\}\cdot\Sigma^{*}\subseteq\{p,q\}\cdot\Sigma^{*}$, hence by Lemma 4.2 we have
$\langle\{t,s\}\rangle\subseteq\langle\{p,q\}\rangle\Rightarrow\langle(t,s)\rangle=\langle(p,q)\rangle$
and thus $\langle\{t,s\}\rangle$ has to be an atom as well. ∎
[…]
  1. 1.
    Construct the pairs-digraph $\mathcal{P}$ whose vertices are the pairs $\binom{Q}{2}$, and where there is an edge $\{p,q\}\rightarrow\{s,t\}$ whenever there exists a letter $a\in\Sigma$ such that $\{p,q\}\cdot a=\{s,t\}$.
  2. 2.
    Compute the strongly connected components of $\mathcal{P}$, and let $\sim$ be the equivalence relation defined by the partition induced by these components. Let $\mathcal{P}/_{\sim}$ be the corresponding quotient graph.
  3. 3.
[…]
Note that the algorithm $\text{Atom}(\mathcal{A})$ can be implemented in polynomial time with respect to the size of the input automaton.
@@
We conclude this section on congruences with the following construction, which will be useful in the sequel. Let $\mathcal{A}$ be a DFA, and let $\sigma\in\text{Cong}(\mathcal{A})$. Given a congruence $\rho\in\text{Cong}(\mathcal{A}/\sigma)$, we define the lifting of $\rho$ as the congruence $\overline{\rho}\in\text{Cong}(\mathcal{A})$ given by:
@@
Consider the mapping $\rho\mapsto\overline{\rho}$, where $\rho\in\text{Cong}(\mathcal{A}/\sigma)$. This defines an injective function
$\text{Cong}(\mathcal{A}/\sigma)\longrightarrow\text{Cong}(\mathcal{A}),$
which is clearly order-preserving. Since the meet and join operations are preserved under this embedding, it follows that $\text{Cong}(\mathcal{A}/\sigma)$ is isomorphic to a sublattice of $\text{Cong}(\mathcal{A})$ via this map. ∎
From the few examples we have considered, it seems that the structure of the congruence poset of a given DFA is somewhat rigid, although we still lack sufficiently complex examples to properly assess the problem. This motivates the following broad open problem and direction of research:
Open Problem 4.7.
Let $P$ be a finite lattice. Does there exist a synchronizing DFA $\mathcal{A}$ such that $\text{Cong}(\mathcal{A})\cong P$ as posets?
@@ 4.1 Semisimplicity vs congruences
By the nihilpotency of $\text{Rad}(\mathcal{A}/\sigma)$ in $\Sigma^{*}/\text{Syn}(\mathcal{A}/\sigma)$, we have that $\exists k\geq 1$ such that:
@@
But then, by the nihilpotency of $\text{Rad}(\mathcal{A})$ in $\Sigma^{*}/\text{Syn}(\mathcal{A})$ we have that $\exists m\geq 1$ such that $\text{Rad}(\mathcal{A})^{m}\subseteq\text{Syn}(\mathcal{A})$, and thus:
@@
and thus we have that $\text{Rad}(\mathcal{A}/\sigma)$ is a nihilpotent ideal of $\Sigma^{*}/\text{Syn}(\mathcal{A})$. Being $\text{Rad}(\mathcal{A})$ the biggest nihilpotent ideal of $\Sigma^{*}/\text{Syn}(\mathcal{A})$, we must have $\text{Rad}(\mathcal{A}/\sigma)\subseteq\text{Rad}(\mathcal{A})$, hence $\text{Rad}(\mathcal{A})=\text{Rad}(\mathcal{A}/\sigma)$. ∎
@@
$\text{Syn}(\mathcal{A}/\tau)\subseteq\text{Syn}(\mathcal{A}/\sigma)\quad\text{and}\quad\text{Rad}(\mathcal{A}/\tau)\subseteq\text{Rad}(\mathcal{A}/\sigma).$
@@
we have that $\mathcal{A}/\sigma$ semisimple.
@@
It is easy to show that $\mathcal{A}/\sigma_{1}\cong\mathcal{C}_{4}$, where $\mathcal{C}_{4}$ stands for the Černý automaton of order $4$ which is simple (see [30]), and $\mathcal{A}/\sigma_{2}$ is a 22-state automaton, and thus it is simple. Observe that $\sigma_{1}\cap\sigma_{2}=\Delta_{\mathcal{A}}$, and by means of Corollary 4.12 we get that $\mathcal{A}$ is semisimple with:
@@ 5 Galois connection
In this section, we establish the existence of a Galois connection between the lattice $\operatorname{Cong}(\mathcal{A})$ of congruences on an automaton $\mathcal{A}$ and the lattice $(\operatorname{Id}(\mathcal{A}),\subseteq)$ of ideals of the transition monoid $M(\mathcal{A})$, where, without loss of generality, we sometimes consider such connections between $\operatorname{Cong}(\mathcal{A})$ and the lattice $(\operatorname{Id}(\mathcal{A}),\subseteq)$ of (two-sided) ideals of the free monoid $\Sigma^{*}$. This is because it is often more convenient to view ideals as languages, that is, as subsets of $\Sigma^{*}$, while working with ideals of $M(\mathcal{A})$ is sometimes preferable due to the finiteness of the monoid, which ensures a finite ideal lattice.
@@
We can extend the above definition on the lattice $(\mathbb{I}\!\operatorname{d}(\mathcal{A}),\subseteq)$ as follows. Let us define
@@
For any $\sigma\in\text{Cong}(\mathcal{A})$ we have $\text{Syn}(\mathcal{A}/\sigma)\cdot I(\sigma)\subseteq\text{Syn}(\mathcal{A})$. Moreover, the projection of the ideal $I(\sigma)\cap\text{Syn}(\mathcal{A}/\sigma)$ into $\mathcal{A}^{*}$ is a nihilpotent ideal of order at most two.
@@
Let $(A,\leq),(B,\leq)$ be posets with $f:A\rightarrow B,\ g:B\rightarrow A$ antimorphisms. We say that the two atimorphisms $f,g$ form a Galois connection if
@@
Let $J\in\operatorname{Id}(\mathcal{A})$ and $u\in J$, then $\text{ker}(u)\supseteq\rho(J)$ which implies $u\in I(\rho(J))$. Hence, we have the inclusion $J\subseteq I(\rho(J))$. Let now consider $\sigma\in\text{Cong}(\mathcal{A})$. Observe that for any $u\in I(\sigma)$ we have $\text{ker}(u)\supseteq\sigma$, from which we deduce $\sigma\subseteq\rho(I(\sigma))$. ∎
We recall a basic fact in Galois connections theory, see [12].
@@
Let $I\subseteq\Sigma^{*}$ be an ideal, $\sigma\in\text{Cong}(\mathcal{A})$, $\sigma:=\rho(I)$ and $J:=I(\sigma)$. Then we have that:
@@
Let $\mathcal{A}$ be a semisimple DFA, and let $I\subseteq\text{M}(\mathcal{A})$ and $\sigma\in\text{Cong}(\mathcal{A})$ be such that $I(\rho(I))=I$ and $\rho(I(\sigma))=\sigma$. Then, for any $u\in\Sigma^{*}$:
$\begin{split}u\cdot I(\sigma)\subseteq\text{Syn}(\mathcal{A})&\Leftrightarrow u\in\text{Syn}(\mathcal{A}/\sigma),\\ u\cdot I\subseteq\text{Syn}(\mathcal{A})&\Leftrightarrow u\in\text{Syn}(\mathcal{A}/\rho(I)).\end{split}$
@@
Assume $u\cdot I(\sigma)\subseteq\text{Syn}(\mathcal{A})$ and suppose, for contradiction, that $u\notin\text{Syn}(\mathcal{A}/\sigma)$. Then there exist $p,q\in Q\cdot u$ such that $p]_{\sigma}\neq[q]_{\sigma}$ and $p\cdot v=q\cdot v$ for every $v\in I(\sigma)$. By definition of $\rho(I(\sigma))$, this implies $(p,q)\in\rho(I(\sigma))=\sigma$, which contradicts $p]_{\sigma}\neq[q]_{\sigma}$. Therefore, $u\in\text{Syn}(\mathcal{A}/\sigma)$. The other implication is a consequence of the definitions.
The second implication is proved similarly. Suppose $u\cdot I\subseteq\text{Syn}(\mathcal{A})$ and $u\notin\text{Syn}(\mathcal{A}/\rho(I))$. Then there exist $p,q\in Q\cdot u$ such that $p]_{\rho(I)}\neq[q]_{\rho(I)}$ and $p\cdot v=q\cdot v$ for every $v\in I$. But then $(p,q)\in\rho(I)$, contradicting the assumption that $p]_{\rho(I)}\neq[q]_{\rho(I)}$. Hence, $u\in\text{Syn}(\mathcal{A}/\rho(I))$. The other implication follows easily from the definitions. ∎
In particular, the previous proposition shows that the left annihilators of $I(\sigma)$ and $I$ seen as ideals of $\mathcal{A}^{*}$ correspond to the projection of $\text{Syn}(\mathcal{A}/\sigma)$ and $\text{Syn}(\mathcal{A}/\rho(I))$, respectively, under the natural projection onto $\mathcal{A}^{*}:=\text{M}(\mathcal{A})/\text{Syn}(\mathcal{A})$.

### v2

@@ 4 The lattice of congruences of an automaton
This section is dedicated to the study of the lattice of congruences of a given DFA $\mathcal{A}=(Q,\Sigma,\delta)$. Using the language of universal algebra, we may view an automaton as a unary algebra on $Q$ with $\Sigma=\{a_{1},\ldots,a_{m}\}$ unary operations, see for instance [8]. Therefore, the poset of congruences $\text{Cong}(\mathcal{A})$, ordered by inclusion, is a finite lattice where for every $\sigma,\tau\in\text{Cong}(\mathcal{A})$ one has
$\sigma\wedge\tau:=\sigma\cap\tau,\qquad\sigma\vee\tau:=\bigcap\{\rho\in\text{Cong}(\mathcal{A})\mid\rho\supseteq\sigma,\tau\}.$
The maximum (top element) is given by the universal relation $1:=\nabla_{\mathcal{A}}=Q\times Q$, and the minimum (bottom element) by the identity relation $0:=\Delta_{\mathcal{A}}=\{(q,q):q\in Q\}$. In lattice-theoretic terms, a non-trivial congruence $\theta\in\text{Cong}(\mathcal{A})$, $\theta\neq 0$, is called an atom if it covers 0, that is, if there is no congruence strictly between 0 and $\theta$. Note that two distinct atoms $\sigma,\tau$ are orthogonal, i.e., $\sigma\cap\tau=0$. If a DFA admits a unique atom, such congruence is called monolith. In other words, a monolith (if exists) is the non-trivial minimum of $\text{Cong}(\mathcal{A})$.
For any pair of states $p,q\in Q$, we define the principal congruence, or the congruence generated by $\{p,q\}$, denoted by $\langle\{p,q\}\rangle$, as follows:
$\langle\{p,q\}\rangle:=\bigcap\left\{\sigma\in\operatorname{Cong}(\mathcal{A})\ \middle|\ (p,q)\in\sigma\right\}.$
Calculating this congruence is quite efficient, and it can be done in $O(m|\mathcal{A}|)$ time, see [16, Theorem 4]. The main strategy of this algorithm is based on the following fact.
@@
With the above notation, $\langle\{p,q\}\rangle$ is equal to the equivalence closure $\langle S\rangle_{eq}$ of the set $S$ defined by:
$S=\{p,q\}\cdot\Sigma^{*}:=\left\{(t,s)\in Q\times Q\ \middle|\ \exists u\in\Sigma^{*}\text{ such that }\{p,q\}\cdot u=\{t,s\}\right\}.$
@@
It follows from [8, Theorem 5.5] applied to a unary algebra. ∎
The following lemma easily follows from the definition of atom of a lattice.
@@
Let $\text{Cong}(\mathcal{A})$ seen as a lattice. Then, for every atom $\sigma\in\text{Cong}(\mathcal{A})$ we have $\sigma=\langle\{p,q\}\rangle$ for some $p,q\in Q$ with $p\neq q$.
For a finite algebra $A$, the Demel-Demlová-Koubek (DDK) algorithm computes the atoms of the congruence lattice of $A$ in $O(m|A|^{r+1})$ time, where $r$ is the maximum arity of the operations in the algebra $A$, and $m$ is the number of operations, see [10, Theorem 4.3]. For a deterministic finite automaton $\mathcal{A}$, the transitions are unary operations ($r=1$), which specializes the complexity to $O(m|Q|^{2})$. For the sake of completeness, since it is also recalled in the computation of the radical ideal, we outline a similar algorithm here. Consider the following preorder on
$\binom{Q}{2}:=\{X\subseteq Q\mid|X|=2\},$
@@
$\{p,q\}]_{\sim}\preceq[\{s,t\}]_{\sim}\quad\text{if and only if}\quad\{p,q\}\preceq\{s,t\}.$
It is straightforward to check that $\{p,q\}\sim\{s,t\}\quad\Longleftrightarrow\quad\{p,q\}\cdot\Sigma^{*}=\{s,t\}\cdot\Sigma^{*}.$ Hence, by Lemma 4.1, if $\{p,q\}\sim\{s,t\}$ then $\langle\{p,q\}\rangle=\langle\{s,t\}\rangle$. Therefore, the map
$L:\binom{Q}{2}/\!\sim\;\longrightarrow\;\text{Cong}(\mathcal{A}),\qquad L([\{p,q\}]_{\sim}):=\langle\{p,q\}\rangle,$
is a well-defined, surjective, antimorphism of posets. Recall that the coatoms of a lattice are its maximal proper elements. It is straightforward to check that the preimage of an atom of $\text{Cong}(\mathcal{A})$ contains a coatom of $\binom{Q}{2}$$_{\sim}$. Hence, the set of atoms of $\text{Cong}(\mathcal{A})$ is in bijection with a subset of the coatoms of $\binom{Q}{2}$$_{\sim}$. The poset $\binom{Q}{2}$$_{\sim}$ admits a natural graph-theoretic interpretation. Consider the pairs digraph $\mathcal{P}$ whose vertex set is $\binom{Q}{2}$, and where there is a directed edge $\{p,q\}\longrightarrow\{s,t\}$ whenever there exists a letter $a\in\Sigma$ such that $\{p,q\}\cdot a=\{s,t\}$. The preorder $\preceq$ coincides with reachability in $\mathcal{P}$, while the equivalence relation $\sim$ corresponds to mutual reachability. Thus, each equivalence class $\{p,q\}]_{\sim}$ is precisely a strongly connected component (SCC) of $\mathcal{P}$ in the usual graph-theoretic sense. Consequently, the Hasse diagram of $\binom{Q}{2}$$_{\sim}$ is isomorphic to the graph whose vertices are the SCC of $\mathcal{P}$ and where one SCC $\{p,q\}]_{\sim}$ is below $\{p^{\prime},q^{\prime}\}]_{\sim}$ if there is a directed edge connecting $\{p,q\}]_{\sim}$ to $\{p^{\prime},q^{\prime}\}]_{\sim}$. In this representation, the coatoms of $\binom{Q}{2}$$_{\sim}$ are in one-to-one correspondence with the sink-SCC, that is, SCCs of $\mathcal{P}$ that have no outgoing edges to other SCCs. This interpretation leads to the following meta-algorithm $\textnormal{Atom}(\mathcal{A})$ for computing the atoms of $\text{Cong}(\mathcal{A})$. The algorithm is a modification of Tarjan’s classical procedure for computing strongly connected components. The pairs digraph $\mathcal{P}$ is explored via a depth-first search. During the execution, the algorithm maintains a global variable $S\subseteq\binom{Q}{2}$ whose role is to record pairs $\{p,q\}$ such that the corresponding equivalence class $\{p,q\}]_{\sim}$ is known not to correspond to an atom, or for which there is another non-trivial pair $(u,v)\in\langle\{p,q\}\rangle$ not in $S$. Whenever the algorithm identifies a sink strongly connected component $C=[\{s,t\}]_{\sim}$, it performs a minimality check by computing the congruence $\langle\{s,t\}\rangle$ and testing whether every non-trivial pair of this congruence already belongs to $S$. If this is the case, then $\langle\{s,t\}\rangle$ is minimal among non-trivial congruences and hence defines an atom of $\text{Con}(\mathcal{A})$. Otherwise, there exists a pair $(u,v)\in\langle\{s,t\}\rangle$ such that $\{u,v\}\notin S$. In this case, one has $\langle\{u,v\}\rangle\subseteq\langle\{s,t\}\rangle$, so $\langle\{u,v\}\rangle$ is a candidate to be an atom. The algorithm then adds all pairs in $C$ to $S$ and recursively applies the same procedure to the SCC $\{u,v\}]_{\sim}$. We emphasize that the main loop of the procedure is executed only for pairs $\{u,v\}\notin S$. Once a pair is inserted into $S$, it is never considered again as a candidate for generating an atom. This guarantees that each pair in $\binom{Q}{2}$ is processed at most once in the minimality checks. As a consequence, the overall cost of the algorithm is $O(m\cdot|Q|^{2})$.
@@
We conclude this section regarding congruences of an automaton with the following construction, which will be useful in the sequel. Let $\mathcal{A}$ be a DFA, and let $\sigma\in\text{Cong}(\mathcal{A})$. Given a congruence $\rho\in\text{Cong}(\mathcal{A}/\sigma)$, we define the lifting of $\rho$ as the congruence $\overline{\rho}\in\text{Cong}(\mathcal{A})$ given by:
@@
By [8, Theorem 6.20] we have that $\text{Cong}(\mathcal{A}/\sigma)\cong[\sigma,\nabla_{\mathcal{A}}]$, where $\sigma,\nabla_{\mathcal{A}}]$ is the sublattice of $\text{Cong}(\mathcal{A})$ defined as follows:
$\sigma,\nabla_{\mathcal{A}}]:=\{\tau\in\text{Cong}(\mathcal{A})\ |\ \sigma\subseteq\tau\}.$
Here the isomorphism is given by $\rho\mapsto\overline{\rho}$. ∎
@@ 4.1 Semisimplicity vs congruences
By the nilpotency of $\text{Rad}(\mathcal{A}/\sigma)$ in $\Sigma^{*}/\text{Syn}(\mathcal{A}/\sigma)$, we have that $\exists k\geq 1$ such that:
@@
But then, by the nilpotency of $\text{Rad}(\mathcal{A})$ in $\Sigma^{*}/\text{Syn}(\mathcal{A})$ we have that $\exists m\geq 1$ such that $\text{Rad}(\mathcal{A})^{m}\subseteq\text{Syn}(\mathcal{A})$, and thus:
@@
and thus we have that $\text{Rad}(\mathcal{A}/\sigma)$ is a nilpotent ideal of $\Sigma^{*}/\text{Syn}(\mathcal{A})$. Being $\text{Rad}(\mathcal{A})$ the biggest nilpotent ideal of $\Sigma^{*}/\text{Syn}(\mathcal{A})$, we must have $\text{Rad}(\mathcal{A}/\sigma)\subseteq\text{Rad}(\mathcal{A})$, hence $\text{Rad}(\mathcal{A})=\text{Rad}(\mathcal{A}/\sigma)$. ∎
@@
$\text{Syn}(\mathcal{A}/\sigma)\subseteq\text{Syn}(\mathcal{A}/\tau)\quad\text{and}\quad\text{Rad}(\mathcal{A}/\sigma)\subseteq\text{Rad}(\mathcal{A}/\tau).$
@@
we have that $\mathcal{A}/\sigma$ is semisimple.
@@
It is easy to show that $\mathcal{A}/\sigma_{1}\cong\mathcal{C}_{4}$, where $\mathcal{C}_{4}$ stands for the Černý automaton of order $4$ which is simple (see [33]), and $\mathcal{A}/\sigma_{2}$ is a 2-state automaton, and thus it is simple. Observe that $\sigma_{1}\cap\sigma_{2}=\Delta_{\mathcal{A}}$, and by means of Corollary 4.8 we get that $\mathcal{A}$ is semisimple with:
@@ 5 Galois connection
In this section, we establish the existence of a Galois connection between the lattice $\operatorname{Cong}(\mathcal{A})$ of congruences on an automaton $\mathcal{A}$ and the lattice $(\operatorname{Id}(\mathcal{A}),\subseteq)$ of ideals of the transition monoid $\text{M}(\mathcal{A})$, where, without loss of generality, we sometimes consider such connections between $\operatorname{Cong}(\mathcal{A})$ and the lattice $(\operatorname{Id}(\mathcal{A}),\subseteq)$ of (two-sided) ideals of the free monoid $\Sigma^{*}$. This is because it is often more convenient to view ideals as languages, that is, as subsets of $\Sigma^{*}$, while working with ideals of $\text{M}(\mathcal{A})$ is sometimes preferable due to the finiteness of the monoid, which ensures a finite ideal lattice.
@@
We can extend the above definition on the poset of two-sided ideals of $\Sigma^{*}$ with respect to the inclusion as follows. Let us define
@@
For any $\sigma\in\text{Cong}(\mathcal{A})$ we have $\text{Syn}(\mathcal{A}/\sigma)\cdot I(\sigma)\subseteq\text{Syn}(\mathcal{A})$. Moreover, the projection of the ideal $I(\sigma)\cap\text{Syn}(\mathcal{A}/\sigma)$ into $\mathcal{A}^{\star}$ is a nilpotent ideal of order at most two.
@@
Let $(A,\leq),(B,\leq)$ be posets with $f:A\rightarrow B,\ g:B\rightarrow A$ antimorphisms. We say that the two antimorphisms $f,g$ form a Galois connection if
@@
Let $J\in\operatorname{Id}(\mathcal{A})$ and $u\in J$, then $\text{ker}(u)\supseteq\rho(J)$ which implies $u\in I(\rho(J))$. Hence, we have the inclusion $J\subseteq I(\rho(J))$. Let us now consider $\sigma\in\text{Cong}(\mathcal{A})$. Observe that for any $u\in I(\sigma)$ we have $\text{ker}(u)\supseteq\sigma$, from which we deduce $\sigma\subseteq\rho(I(\sigma))$. ∎
Let us now state the following basic result concerning Galois connections:
@@
Let $K\subseteq\Sigma^{*}$ be an ideal, $\tau\in\text{Cong}(\mathcal{A})$, $\sigma:=\rho(K)$ and $J:=I(\tau)$. Then we have that:
@@
Let $\mathcal{A}$ be a semisimple DFA, and let $K\in\operatorname{Id}(\mathcal{A})$ and $\tau\in\text{Cong}(\mathcal{A})$ be such that $I(\rho(K))=K$ and $\rho(I(\tau))=\tau$. Then, for any $u\in\Sigma^{*}$:
$\begin{split}u\cdot I(\tau)\subseteq\text{Syn}(\mathcal{A})&\Leftrightarrow u\in\text{Syn}(\mathcal{A}/\tau),\\ u\cdot K\subseteq\text{Syn}(\mathcal{A})&\Leftrightarrow u\in\text{Syn}(\mathcal{A}/\rho(K)).\end{split}$
@@
Assume $u\cdot I(\tau)\subseteq\text{Syn}(\mathcal{A})$ and suppose, for contradiction, that $u\notin\text{Syn}(\mathcal{A}/\tau)$. Then there exist $p,q\in Q\cdot u$ such that $p]_{\tau}\neq[q]_{\tau}$ and $p\cdot v=q\cdot v$ for every $v\in I(\tau)$. By definition of $\rho(I(\tau))$, this implies $(p,q)\in\rho(I(\tau))=\tau$, which contradicts $p]_{\tau}\neq[q]_{\tau}$. Therefore, $u\in\text{Syn}(\mathcal{A}/\tau)$. The other implication is an immediate consequence of Lemma 5.4.
The second implication is proved similarly: suppose $u\cdot K\subseteq\text{Syn}(\mathcal{A})$ and $u\notin\text{Syn}(\mathcal{A}/\rho(K))$. Then there exist $p,q\in Q\cdot u$ such that $p]_{\rho(K)}\neq[q]_{\rho(K)}$ and $p\cdot v=q\cdot v$ for every $v\in K$. But then $(p,q)\in\rho(K)$, contradicting the assumption that $p]_{\rho(K)}\neq[q]_{\rho(K)}$. Hence, $u\in\text{Syn}(\mathcal{A}/\rho(K))$. The other implication follows easily combining Lemma 5.4 and Proposition 5.6. ∎
In particular, the previous proposition shows that the left annihilators of $I(\tau)$ and $K$ seen as ideals of $\mathcal{A}^{\star}$ correspond to the projection of $\text{Syn}(\mathcal{A}/\tau)$ and $\text{Syn}(\mathcal{A}/\rho(K))$, respectively, under the natural projection onto $\mathcal{A}^{\star}=\text{M}(\mathcal{A})/\text{Syn}(\mathcal{A})$.